OpenStudy (anonymous):

20) The Everton College store paid $1804 for an order of 48 calculators. The store paid $9 for each scientific calculator. The others all graphing calculators cost $58 each. How many of each type of calculator was ordered?

OpenStudy (amistre64):

c@9 = (1804 - 48(9))/(58-9) maybe .... i always get this one confused

OpenStudy (anonymous):

solve \[9x+58(84-x)=1804\] so get your answer

OpenStudy (anonymous):

sensei always does it in one step. probably right

OpenStudy (anonymous):

anyway what i wrote was wrong. solve \[9x+58(48-x)=1804\]

OpenStudy (amistre64):

lets see if im right :) 28(9) + 20(58) ?= 1804 252 + 1160 ?= 1804 1412!=! 1804 ... might got em backwards then ...

OpenStudy (amistre64):

yep; had em backwards lol

OpenStudy (anonymous):

reasoning as follows. x is number of $9 for a total of 9x dollars. if x are $9 calculators then the rest (48-x) cost $58 for a total of 58(48-x) dollars.

OpenStudy (anonymous):

therefore one expression for the total is \[9x+58(48-x)\] and the other expression for the total is $1084 because you are told that

OpenStudy (anonymous):

so set \[9x+58(48-x)=1084\] and solve for x

OpenStudy (amistre64):

1804 .. but yeah

OpenStudy (amistre64):

b@y = (z -t(x))/(y-x) amount@58 = (1804 - 48(9))/(58-9)